OpenStudy (anonymous):

: Simplify the expression below. 7x(6x) Select one of the options below as your answer:A. 13xB. 13x2C. 42xD. 42x2

OpenStudy (anonymous):

try that last method I showed you on this one.

OpenStudy (anonymous):

7*6?

OpenStudy (agent0smith):

Multiply the numbers and the x's together.. might help to re-write it like this: \[7 \times 6 \times x \times x\]

OpenStudy (anonymous):

yeah, that part first

OpenStudy (agent0smith):

\[(7*6) (x* x)\]